OpenStudy (anonymous):

Use the integral test to determine whether the series En=1- infinity of 1/(3n +1) Sorry I don't have the latex script to properly type this on my iPad. Please let me know if the question doesn't make sense

OpenStudy (anonymous):

U will need to find this integral: \[\int\limits_{1}^{\infty} \frac{ 1 }{ 3n+1 }\] If this integral converges, your series converges too. The only criteria is that the function is positive valued, continuous and decreasing. Check, check, check

OpenStudy (anonymous):

That integral is , 1/3 ln (3n+1) +C

OpenStudy (anonymous):

As \[\lim_{n \rightarrow \infty} \ln n =\infty \] thus your series is divergent.

OpenStudy (anonymous):

Thank you, Andrus!!

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

*Andras

OpenStudy (anonymous):

The answer is correct, just as a side note: Make sure that in an exam you first verify that the sequence \(a_n\geq 0, \ \forall n \in \mathbb{N}\) and that the series is monotone decreasing \(a_{n+1} \leq a_n, \ \forall n \in \mathbb{N}\) otherwise you can't apply the integral test.

OpenStudy (anonymous):

@Andras has already mentioned that in his post, just to give those lines credits here, try \(f'(x) \leq 0\) by transforming the sequence into a real valued function \(\mathbb{R} \to \mathbb{R}\)

OpenStudy (anonymous):

I did :) It was the check, check, check part

OpenStudy (anonymous):

Thank you both!!